OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received: x − 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

OpenStudy (mathstudent55):

You are looking for the number of votes Maya received. You are told that the number of votes Maya received is x. That means in the equation you were given you are looking for x. y in the equation is the total number of votes cast. You are told that y = 125,000. Substitute y with 125,000 in the equation, and solve for x.

OpenStudy (mathstudent55):

x - 62,500 = 0.25y Replace y with 125,000. You get: x - 62,500 = 0.25 * 125,000 Now do the multiplication on the right side: x - 62,500 = 31,250 Can you solve for x now? Hint: Add 62,500 to both sides.

OpenStudy (mathstudent55):

What did you get?

OpenStudy (anonymous):

9375

OpenStudy (mathstudent55):

Hoe can a number in the 30 thousands plus a number in the 60 thousands equal less than 10 thousand? It has to be close to 100 thousand. You are missing a zero at the end.

OpenStudy (anonymous):

so is that the answer tho

OpenStudy (mathstudent55):

The answer is 93750